# thanks: corresponding author

### v3

@@ II Topological surgery
Figure 1: The $3$-dimensional 0-surgery removes an embedding of $S^{0}\times D^{3}$ and replaces it with $S^{1}\times S^{2}$. This process, around the critical point, can be represented by a Morse function such as $f=-x^{2}+y^{2}+z^{2}$ (suppressing one dimension), whose level sets are shown above.
@@ III.2 The kink number
$\operatorname{kink}(\partial W;V^{\mu})=\sum_{j=1}^{m}\penalty 10000\ \operatorname{kink}(\Sigma_{j};V^{\mu}).$
@@ III.3 Trivial 2+1 cobordisms
where $g(\theta)$ varies smoothly and monotonically from 0 to $1$ as $\theta$ goes from $\theta_{-}$ to $\theta_{+}$, with $g^{\prime}(\theta_{\pm})=0$. The square-root factor in the denominator ensures that the vector field has unit norm. Both the functional form of $g(\theta)$ and the interval $\Delta\theta\equiv\theta_{+}-\theta_{-}$ are arbitrary; the integral is independent of these choices. We then find [^4] Recall that, by convention, the outgoing normal is negative, so $f^{t}$ is negative in this region.
@@
For the top strip, the calculation is analogous; here the vector field changes from tangent to outward-normal:
@@ III.4 The wormhole nucleation cobordism
$\chi(\Sigma_{i})=2=\frac{1}{2}\chi(\partial\Sigma_{i})=1-g.$
Hence we must have g=1g=-1, which is impossible. Thus, the condition for an everywhere non-singular Lorentzian metric cannot be met for this simple cobordism; a singularity of index 1-1 is required [^5] Since the initial manifold is assumed to be simply connected, we take $g=0$. Thus, from the Poincaré-Hopf theorem, the index is 1g=1-1-g=-1. If the genus was positive, the index would be lower, but since the gradient of a Morse function can only have singularities of index $\pm 1$, that would correspond to multiple singularities of index 1-1 and therefore multiple successive surgeries.—consistent with the index of the Morse gradient for a 0-surgery, cf. (1).
@@ III.5 Spin structure
$\operatorname{rank}(h)\bmod 2=0\quad\Longleftrightarrow\quad w_{2}(M)=0.$
@@
$\text{Spin}^{c}(4)\cong SL(2,\mathbb{C})\times_{\mathbb{Z}_{2}}U(1).$
@@ IV.2 The Morse spacetime
The choice of the Riemannian metric is arbitrary. Here we consider the level sets of the Morse function as embedded in $\mathbb{R}^{5}$, so that a natural Riemannian metric is the one induced on the level sets from the Euclidean metric in $\mathbb{R}^{5}$:
@@
with $\mathrm{d}\Omega_{2}^{2}$ the volume element on $S^{2}$ and $U\equiv U(z,r)=r^{2}+z^{2}+1$. Note that the metric has a curvature singularity at the critical point $z=r=0$ of the Morse function, as expected from the asymptotic behavior of the curvature tensor near the critical point [28, 29]:
@@
To analyze the behavior of geodesics in this spacetime we note that it possesses three Killing vectors, arising from the spherical symmetry; they are preserved by the Lie bracket between $\partial_{\mu}f$ and the Killing vectors of $g^{R}_{\mu\nu}$ (see Theorem IV.2):
@@
We conclude this section with a brief analysis of the violations of the energy conditions. The analysis is conducted in the coordinates $\{z,r,\theta,\phi\}$, in which the metric takes the form (13). This yields a slight simplification in the expression that classifies the Hawking–Ellis types of the stress–energy tensor. By defining the functions
$C$ $\equiv C(z,r)=z^{2}r^{2}+z^{4}-r^{2}+z^{2},$
$F$ $\equiv F(z,r)=z^{2}+r^{2},$
a tetrad for the metric (13) is given by
$\mathbf{e}_{z}$ $=-\sqrt{\frac{C}{F}}\,\mathrm{d}z+\frac{zr(F+2)}{\sqrt{CF}}\,\mathrm{d}r,$ (18a)
$\mathbf{e}_{r}$ $=\sqrt{\frac{F(F+1)}{C}}\,\mathrm{d}r,$ (18b)
$\mathbf{e}_{\theta}$ $=r\,\mathrm{d}\theta,\qquad\mathbf{e}_{\phi}=r\sin(\theta)\,\mathrm{d}\phi.$ (18c)
To classify the stress-energy tensor into the different Hawking-Ellis types, one needs to investigate its eigenvalues, as shown in Refs. [58, 82, 83]. In many cases, however, as pointed out in Ref. [83], the starting tetrad is not the canonical one in which the stress-energy tensor type is explicit, and a Lorentz transformation may be needed to bring it to canonical form.
@@
Figure 9: The stress–energy tensor changes type in different regions of spacetime. The pastel-red regions correspond to type I, while the light-blue regions correspond to type IV. At the boundary the tensor is type II. The graph also depicts the level sets of the Morse function in $(t,r)$ coordinates, revealing that the energy‑condition violations associated with type-IV regions occur on the critical level set.
@@ IV.3 A singular Lorentzian metric on the complex projective plane
$U_{2}:\quad-\left(\zeta_{2}\zeta_{1}+1\right)\partial_{\zeta_{1}}-\zeta_{2}^{2}\partial_{\zeta_{2}},$
$U_{3}:\quad-\xi_{2}\partial_{\xi_{1}}+\partial_{\xi_{2}},$
@@
$w^{\mu}=\begin{cases}{\begin{aligned} U_{1}:(x^{2}-y^{2}+z)\,\partial_{x}+(2xy+t)\,\partial_{y}+(xz-ty)\,\partial_{z}+(tx+yz)\,\partial_{t},\end{aligned}}\\ {\begin{aligned} U_{2}:-(ac-bd+1)\,\partial_{a}-(ad+bc)\,\partial_{b}-(c^{2}-d^{2})\,\partial_{c}-2cd\,\partial_{d},\end{aligned}}\\ U_{3}:{-\gamma\,\partial_{\alpha}-\delta\,\partial_{\beta}+\partial_{\gamma},}\end{cases}$ (21)
@@
where $g^{R}_{\mu\nu}$ is a Riemannian metric and $V_{\mu}$ is a nowhere-vanishing line field obeying $g_{R}^{\mu\nu}V_{\mu}V_{\nu}=1$. Then $\mathcal{L}_{\xi}g^{L}_{\mu\nu}=0$ iff
@@
$\mathcal{L}_{\xi}g^{R}_{\mu\nu}=2\left[(\mathcal{L}_{\xi}V_{\mu})V_{\nu}+V_{\mu}(\mathcal{L}_{\xi}V_{\nu})\right].$
Taking the Lie derivative of the normalization condition $g_{R}^{\mu\nu}V_{\mu}V_{\nu}=1$ and using $\mathcal{L}_{\xi}g^{L}_{\mu\nu}=0$ yields $(\mathcal{L}_{\xi}V_{\mu})V_{\nu}=0$. Because $V_{\mu}$ is nowhere zero, this implies $\mathcal{L}_{\xi}V_{\mu}=0$, which in turn forces $\mathcal{L}_{\xi}g^{R}_{\mu\nu}=0$. ∎
By Theorem IV.2, the Lorentzian metric (22) has no Killing vectors, since the Lie bracket of $w_{\mu}$ with each Killing vector of the Fubini–Study metric is non-zero.
@@
The behavior of these timelike geodesics in our construction is not surprising. Since the vector field on the closed manifold $\mathbb{CP}^{2}$ possesses a unique singular point, every maximal integral curve originates and terminates at this singularity. These integral curves determine the future orientation of the light-cones for the metric (22), which explains why the geodesics are drawn toward the singularity. Moreover, apart from the singular point where the aperture vanishes, the light-cones maintain a non-zero opening angle. This permits the formation of CTCs and, accordingly, the emergence of chronological horizons. By the argument in Ref. [32], these horizons must be compactly generated Cauchy horizons.
@@
For example, after numerically computing a tetrad $e_{a}^{\mu}$ for the metric (22), we consider the event $p=(1,0,0,0)$ in the $U_{1}$ chart. With our specific numerical choice of tetrads, we found that the stress-energy tensor $T^{ab}$ at $p$ has complex eigenvalues, with corresponding complex eigenvectors. So $T^{ab}$ should be of Hawking-Ellis type IV [58, 82, 83]. Because the classification varies continuously, an extended region must exist in which the tensor remains of type IV; in that region all the standard energy conditions are violated.
@@ V.2 Interface between the complex projective plane and the annulus
$0\leq\psi\leq 4\pi,\qquad 0\leq\chi\leq\tfrac{\pi}{2}\sqrt{\tfrac{6}{\Lambda}}.$
@@
The last step is to verify the presence of thin shells or distributional terms in the equations of motions. We observe that if the derivatives of the metric are continuous across $\Sigma$,
@@
then $\left[\Gamma^{\alpha}_{\mu\nu}\right]=0$ and condition (27) is satisfied. Equation (32) can be recast as two separate conditions on the jumps of the derivatives of both the Riemannian metric and the line field:
@@
$\dot{a}(\tau_{1})=\frac{1}{2}\cos\left(\chi_{1}\sqrt{\tfrac{\Lambda}{6}}\right),\qquad\dot{b}(\tau_{1})=\cos\left(\chi_{1}\sqrt{\tfrac{\Lambda}{6}}\right)$
where a dot denotes differentiation with respect to $\tau$. These conditions remove any discontinuity in the derivatives of the line field as well, thanks to the preliminary equality $\chi_{1}=\tau_{1}$. Consequently, no distributional terms appear in the equations of motion.
@@ V.3 Interface between the annulus and the Morse spacetime
$\mathrm{d}s^{2}_{\mathrm{M}}=\mathrm{d}s^{2}_{\mathrm{Flat}}+\left(\cos(2\psi)\,\mathrm{d}\xi-\xi\sin(2\psi)\,\mathrm{d}\psi\right)^{2},$ (34)
@@
with $\mathrm{d}\Omega_{2}^{2}$ the volume element on the unit 2-sphere. The Lorentzian metric (12) is then given in compact form as
@@
$\alpha(\sigma_{1})=\xi_{1},\qquad\beta(\sigma_{1})=1.$ (37)
@@
again identified on $\Sigma$. They satisfy $l^{\mu}N_{\mu}\neq 0$ everywhere on $\Sigma$ and have the property that $l^{+\mu}$ is directed into $\mathbb{M}_{+}$, while $l^{-\mu}$ is directed out of $S^{3}_{I}\times I$, thereby satisfying the second preliminary gluing condition (26). To avoid thin shells we follow the procedure of the previous subsection, which leads to the requirements
$\dot{\alpha}(\sigma_{1})=1,\qquad\dot{\beta}(\sigma_{1})=0.$ (38)
@@ V.4 Homotopy between the vector fields along the neck
$a(\tau_{0})=\alpha(\sigma_{0})=R,\qquad\dot{a}(\tau_{0})=\dot{\alpha}(\sigma_{0})=0.$ (40)
@@
$H_{2}(\eta,x^{\mu})=\tfrac{(z-x^{2}-y^{2})\,\partial_{x}+t\,\partial_{y}-(ty+xz^{\,c(\eta)})\,\partial_{z}-(tx-yz^{\,c(\eta)})\,\partial_{t}}{\bigl\lVert(z-x^{2}-y^{2})\,\partial_{x}+t\,\partial_{y}-(ty+xz^{\,c(\eta)})\,\partial_{z}-(tx-yz^{\,c(\eta)})\,\partial_{t}\bigr\rVert},$
@@
$H_{3}(\eta,x^{\mu})=\frac{s_{1}^{\mu}+\bigl(1-b(\eta)\bigr)s_{2}^{\mu}}{\bigl\lVert s_{1}^{\mu}+\bigl(1-b(\eta)\bigr)s_{2}^{\mu}\bigr\rVert},$
so that $H_{3}(\eta_{1/2},x^{\mu})=s_{1}^{\mu}/\lVert s_{1}^{\mu}\rVert$. Finally, $s_{1}^{\mu}$ is related to $v^{\mu}$ by successive rotations in the $xz$, ytyt, and $xy$ planes, which we summarize by a one-parameter rotation matrix $R(\alpha(\eta))$ with $\alpha(\eta_{1/2})=0$ and $\alpha(\eta_{1})=\pi/2$.

### v4

@@ II Topological surgery
Figure 1: The $3$-dimensional 0-surgery removes an embedding of $S^{0}\times D^{3}$ and replaces it with $D^{1}\times S^{2}$. This process, around the critical point, can be represented by a Morse function such as $f=-x^{2}+y^{2}+z^{2}$ (suppressing one dimension), whose level sets are shown above.
@@ III.2 The kink number
$\operatorname{kink}(\partial W;V^{\mu})=\sum_{j=0}^{m}\penalty 10000\ \operatorname{kink}(\Sigma_{j};V^{\mu}).$
@@ III.3 Trivial 2+1 cobordisms
where $g(\theta)$ varies smoothly and monotonically from 0 to $1$ as $\theta$ goes from $\theta_{-}$ to $\theta_{+}$, with $g^{\prime}(\theta_{\pm})=0$. The square-root factor in the denominator ensures that the vector field has unit norm. Both the functional form of $g(\theta)$ and the interval $\Delta\theta\equiv\theta_{+}-\theta_{-}$ are arbitrary; the integral is independent of these choices. We then find
@@
For the top strip, the calculation is analogous; here the vector field changes from tangent to outward-normal [^4] Recall that, by convention, the outgoing normal is negative, so $f^{t}$ is negative in this region.:
@@ III.4 The wormhole nucleation cobordism
$\chi(\Sigma_{i})=\frac{1}{2}\chi(\partial\Sigma_{i})=1-g.$
Hence from $\chi(W)=1$ we must have g=1g=-1, which is impossible. Thus, the condition for an everywhere non-singular Lorentzian metric cannot be met for this simple cobordism; a singularity of index 1-1 is required [^5] Since the initial manifold is assumed to be simply connected, we take $g=0$. Thus, from the Poincaré-Hopf theorem, the index is 1g=1-1-g=-1. If the genus was positive, the index would be lower, but since the gradient of a Morse function can only have singularities of index $\pm 1$, that would correspond to multiple singularities of index 1-1 and therefore multiple successive surgeries.—consistent with the index of the Morse gradient for a 0-surgery, cf. (1).
@@ III.5 Spin structure
$\operatorname{rank}(h)\bmod 2=0\quad\Longleftrightarrow\quad w_{2}(W)=0.$
@@
$\text{Spin}^{c}(1,3)\cong SL(2,\mathbb{C})\times_{\mathbb{Z}_{2}}U(1).$
@@ IV.2 The Morse spacetime
The choice of the Riemannian metric is arbitrary. Here we consider the level sets of the Morse function as embedded in $\mathbb{R}^{5}$, so that a natural Riemannian metric is the one induced on the graph of $f$ from the Euclidean metric in $\mathbb{R}^{5}$:
@@
with $\mathrm{d}\Omega_{2}^{2}$ the line element on $S^{2}$ and $U\equiv U(z,r)=r^{2}+z^{2}+1$. Note that the metric has a curvature singularity at the critical point $z=r=0$ of the Morse function, as expected from the asymptotic behavior of the curvature tensor near the critical point [50, 49]:
@@
To analyze the behavior of geodesics in this spacetime we note that it possesses three Killing vectors, arising from the spherical symmetry; they are preserved by the Lie bracket between $\partial^{\mu}f$ and the Killing vectors of $g^{R}_{\mu\nu}$ (see Theorem IV.2):
@@
We conclude this section with a brief analysis of the violations of the energy conditions. The analysis is conducted in the coordinates $\{z,r,\theta,\phi\}$, in which the metric takes the form (13). This yields a slight simplification in the expression that classifies the Hawking–Ellis types of the stress–energy tensor. By defining the function $F\equiv F(z,r)=z^{2}+r^{2}$ a tetrad for the metric (13) is given by
$\mathbf{e}_{0}$ $=\sqrt{\frac{F+1}{F}}(z\mathrm{d}z-r\mathrm{d}r),$ (18a)
$\mathbf{e}_{1}$ $=\frac{r\mathrm{d}z+z\mathrm{d}r}{\sqrt{F}},$ (18b)
$\mathbf{e}_{2}$ $=r\,\mathrm{d}\theta,\qquad\mathbf{e}_{3}=r\sin(\theta)\,\mathrm{d}\phi.$ (18c)
To classify the stress-energy tensor into the different Hawking-Ellis types, one needs to investigate its eigenvalues, as shown in Refs. [38, Martín–Moruno2017, 58]. In many cases, however, as pointed out in Ref. [58], the starting tetrad is not the canonical one in which the stress-energy tensor type is explicit, and a Lorentz transformation may be needed to bring it to canonical form.
@@
Figure 9: The stress–energy tensor changes type in different regions of spacetime. The pastel-red regions correspond to type I, while the light-blue regions correspond to type IV. At the boundary the tensor is type II. The graph also depicts the level sets of the Morse function in $(z,r)$ coordinates, revealing that the energy‑condition violations associated with type-IV regions occur on the critical level set.
@@ IV.3 A singular Lorentzian metric on the complex projective plane
$U_{2}:\quad-\zeta_{1}^{2}\partial_{\zeta_{1}}-(1+\zeta_{1}\zeta_{2})\partial_{\zeta_{2}},$
$U_{3}:\quad\partial_{\xi_{1}}-\xi_{1}\partial_{\xi_{2}},$
@@
$w^{\mu}=\begin{cases}{\begin{aligned} U_{1}:(x^{2}-y^{2}+z)\,\partial_{x}+(2xy+t)\,\partial_{y}+(xz-ty)\,\partial_{z}+(tx+yz)\,\partial_{t},\end{aligned}}\\ {\begin{aligned} U_{2}:-(a^{2}-b^{2})\,\partial_{a}-2ab\,\partial_{b}-(ac-bd+1)\,\partial_{c}-(ad+bc)\,\partial_{d},\end{aligned}}\\ U_{3}:{\partial_{\alpha}-\alpha\,\partial_{\gamma}-\beta\,\partial_{\delta},}\end{cases}$ (21)
@@
where $g^{R}_{\mu\nu}$ is a Riemannian metric and $V_{\mu}$ is a nowhere-vanishing line field obeying $g_{R}^{\mu\nu}V_{\mu}V_{\nu}=1$. Let $\xi^{\mu}$ be a vector field such that $\mathcal{L}_{\xi}V_{\mu}=\alpha V_{\mu}$, for some smooth function $\alpha$. Then $\mathcal{L}_{\xi}g^{L}_{\mu\nu}=0$ iff
@@
$\mathcal{L}_{\xi}g^{R}_{\mu\nu}=2\left[(\mathcal{L}_{\xi}V_{\mu})V_{\nu}+V_{\mu}(\mathcal{L}_{\xi}V_{\nu})\right]=4\alpha V_{\mu}V_{\nu}.$
Taking the Lie derivative of the normalization condition:
$0=\mathcal{L}_{\xi}(g_{R}^{\mu\nu}V_{\mu}V_{\nu})$ $=(\mathcal{L}_{\xi}g_{R}^{\mu\nu})V_{\mu}V_{\nu}+2g_{R}^{\mu\nu}(\mathcal{L}_{\xi}V_{\mu})V_{\nu}$
$=-4\alpha+2\alpha=-2\alpha$
This implies $\alpha=0$ and so $\mathcal{L}_{\xi}V_{\mu}=0$ and $\mathcal{L}_{\xi}g^{R}_{\mu\nu}=0$. ∎
By Theorem IV.2, the Lorentzian metric (22) has no obvious Killing vectors, since the Lie bracket of $w^{\mu}$ with each Killing vector of the Fubini–Study metric is non-zero. Indeed, it is possible to prove that (22) has no Killing vectors [^7] In the $U_{3}$ chart, if there exists $\xi^{\mu}$ Killing, then $\mathcal{L}_{\xi}R_{\mu\nu\rho\sigma}=0$ implies, at the origin $\bm{0}$, $\left.\xi^{\mu}\right|_{\bm{0}}=0$ and $\left.(\nabla_{\mu}\xi_{\nu})\right|_{\bm{0}}=0$. It follows then that $\xi^{\mu}$ is trivial [89, p. 442].
@@
The behavior of these timelike geodesics in our construction is not surprising. Since the vector field on the closed manifold $\mathbb{CP}^{2}$ possesses a unique singular point, every maximal integral curve originates and terminates at this singularity [^8] This is not generally true, since closed orbits can exist. However, for the vector field $w^{\mu}$, there are no closed orbits; they all begin and end at the field’s zero point in the $U_{1}$ chart.. These integral curves determine the future orientation of the light-cones for the metric (22), which explains why the geodesics are drawn toward the singularity. Moreover, apart from the singular point where the aperture vanishes, the light-cones maintain a non-zero opening angle. This permits the formation of CTCs [^9] An explicit future-directed CTC is given for example by the curve $\gamma^{\mu}(s)=\tfrac{1}{100}\Big(2\sin(s),-2\cos(s),\cos(s),\sin(s)\Big),\ s\in[0,2\pi].$ and, accordingly, the emergence of chronological horizons. By the argument in Ref. [39], these horizons must be compactly generated Cauchy horizons.
@@
For example, after numerically computing a tetrad $e_{a}^{\mu}$ for the metric (22), we consider the event $p=(1,0,0,0)$ in the $U_{1}$ chart. With our specific numerical choice of tetrads, we found that the stress-energy tensor $T^{ab}$ at $p$ has complex eigenvalues, with corresponding complex eigenvectors. So $T^{ab}$ should be of Hawking-Ellis type IV [38, Martín–Moruno2017, 58]. Because the classification varies continuously, an extended region must exist in which the tensor remains of type IV; in that region all the standard energy conditions are violated.
@@ V.2 Interface between the complex projective plane and the annulus
$0\leq\psi\leq 4\pi,\qquad 0\leq\chi<\tfrac{\pi}{2}\sqrt{\tfrac{6}{\Lambda}}.$
@@
The last step is to impose a sufficient condition for the absence of thin shells or distributional terms in the equations of motion. We observe that if the derivatives of the metric are continuous across $\Sigma$,
@@
then $\left[\Gamma^{\alpha}_{\mu\nu}\right]=0$ and condition (27) is satisfied. A sufficient set of matching conditions that ensure the validity of Equation (32) are two separate conditions on the jumps of the derivatives of both the Riemannian metric and the line field:
@@
$\dot{a}(\tau_{1})=\frac{1}{2}\cos\left(\chi_{1}\sqrt{\tfrac{\Lambda}{6}}\right),\qquad\dot{b}(\tau_{1})=\sqrt{\frac{\Lambda}{6}}\cos\left(\chi_{1}\sqrt{\tfrac{\Lambda}{6}}\right)$
where a dot denotes differentiation with respect to $\tau$. These conditions remove any discontinuity in the derivatives of the line field as well, thanks to the preliminary equality $\chi_{1}=\tau_{1}$ and the fact that $k^{+}_{\mu}$ and $k^{-}_{\mu}$ agree to first order across $\Sigma$. Consequently, under this stronger matching conditions the metric is $C^{1}$ across $\Sigma$ and no distributional terms appear in the equations of motion.
@@ V.3 Interface between the annulus and the Morse spacetime
$\mathrm{d}s^{2}_{\mathrm{M}}=\mathrm{d}s^{2}_{\mathrm{Flat}}+\xi^{2}\left(\cos(2\psi)\,\mathrm{d}\xi-\xi\sin(2\psi)\,\mathrm{d}\psi\right)^{2},$ (34)
@@
with $\mathrm{d}\Omega_{2}^{2}$ the line element on the unit 2-sphere. The Lorentzian metric (12) is then given in compact form as
@@
$\alpha(\sigma_{1})=\xi_{1},\qquad\beta(\sigma_{1})=\xi_{1}.$ (37)
@@
again identified on $\Sigma$. They satisfy $l^{\mu}N_{\mu}\neq 0$ everywhere on $\Sigma$ and have the property that $l^{+\mu}$ is directed into $\mathbb{M}_{+}$, while $l^{-\mu}$ is directed out of $S^{3}_{I}\times I$, thereby satisfying the second preliminary gluing condition (26). Applying the same sufficient $C^{1}$-matching conditions as in the previous subsection leads to
$\dot{\alpha}(\sigma_{1})=1,\qquad\dot{\beta}(\sigma_{1})=1.$ (38)
@@ V.4 Homotopy between the vector fields along the neck
$2a(\tau_{0})=\alpha(\sigma_{0})=R,\qquad\dot{a}(\tau_{0})=\dot{\alpha}(\sigma_{0})=0.$ (40)
@@
$H_{2}(\eta,x^{\mu})=\tfrac{(z-x^{2}-y^{2})\partial_{x}+t\partial_{y}-t(y\partial_{z}+x\partial_{t})-(c(z-1)+1)(x\partial_{z}-y\partial_{t})}{\bigl\lVert(z-x^{2}-y^{2})\partial_{x}+t\partial_{y}-t(y\partial_{z}+x\partial_{t})-(c(z-1)+1)(x\partial_{z}-y\partial_{t})\bigr\rVert},$
@@
$H_{3}(\eta,x^{\mu})=\frac{s_{1}^{\mu}+\bigl(1-b(\eta)\bigr)s_{2}^{\mu}}{\bigl\lVert s_{1}^{\mu}+\bigl(1-b(\eta)\bigr)s_{2}^{\mu}\bigr\rVert},\quad\eta\in[\eta_{0},\eta_{1/2}],$
so that $H_{3}(\eta_{1/2},x^{\mu})=s_{1}^{\mu}/\lVert s_{1}^{\mu}\rVert$. Finally, $s_{1}^{\mu}$ is related to $v^{\mu}$ by successive rotations in the $xz$, ytyt, and $xy$ planes:
$R=\begin{pmatrix}0&0&1&0\\ 0&0&0&1\\ 1&0&0&0\\ 0&1&0&0\end{pmatrix}\in SO(4),\quad Rs_{1}=v.$
Since $SO(4)$ is connected, we choose a smooth path $R(\alpha(\eta))$ with $R(\alpha(\eta_{1/2}))=\mathbb{I}$ and $R(\alpha(\eta_{1}))=R$ and set
$H_{4}(\eta,x^{\mu})=R(\alpha(\eta))\penalty 10000\ \frac{s_{1}^{\mu}}{\bigl\lVert s_{1}^{\mu}\bigr\rVert},\quad\eta\in[\eta_{1/2},\eta_{1}].$